# GenExam: A Multidisciplinary Text-to-Image Exam

### v2

@@ Abstract
Exams are a fundamental test of expert-level intelligence and require integrated understanding, reasoning, and generation. Existing exam-style benchmarks mainly focus on understanding and reasoning tasks, and current generation benchmarks emphasize the illustration of world knowledge and visual concepts, neglecting the evaluation of rigorous drawing exams. We introduce GenExam, the first benchmark for multidisciplinary text-to-image exams, featuring 1,000 samples across 10 subjects with exam-style prompts organized under a four-level taxonomy. Each problem is equipped with ground-truth images and fine-grained scoring points to enable a precise evaluation of semantic correctness and visual plausibility. Experiments show that even state-of-the-art models such as GPT-Image-1 and Gemini-2.5-Flash-Image achieve less than 15% strict scores, and most models yield almost 0%, suggesting the great challenge of GenExam. By framing image generation as an exam, GenExam offers a rigorous assessment of models’ ability to integrate understanding, reasoning, and generation, providing insights on the path to general AGI. Our benchmark and evaluation code are released at https://github.com/OpenGVLab/GenExam.
Figure 1: Examples and performance of state-of-the-art text-to-image models on GenExam. Orange dashed circles indicate scoring points. GenExam contains complex and diverse prompts resembling human exams, and pose great challenge to existing models.
@@ 1 Introduction
Exams are the ultimate test of expert-level intelligence. They are not merely about recalling knowledge points, but serve as a comprehensive assessment of understanding, reasoning, and generation. The ability to solve multidisciplinary exam problems indicates that a model has expert-level intelligence surpassing that of most adults. Therefore, exam-style benchmarks naturally serve as a crucial yardstick to evaluate the progress of artificial general intelligence (AGI) towards domain expertise Morris et al. (2023); Yue et al. (2024a); Phan et al. (2025).
However, existing multidisciplinary benchmarks are primarily focused on understanding tasks, including pure text benchmarks Wang et al. (2024b); Zhong et al. (2023) and multimodal benchmarks Yue et al. (2024a; b), but rarely focus on image generation tasks. In the field of text-to-image generation (T2I), current benchmarks focus more on general world knowledge reasoning Niu et al. (2025); Sun et al. (2025); Fu et al. (2024). Some benchmarks have explored the domain of subject-specific images but are largely limited to concept illustration Luo et al. (2025b); Chang et al. (2025b; a), a relatively simple task with loose diagnostic criteria, similar to “illustrating concepts through image generation” rather than “solving a drawing exam”.
Upon a closer inspection of graph-drawing questions in common multidisciplinary exams, such as AP CollegeBoard (1952), A-level Cambridge (1951), and IB IBO (1968), we find that text-to-image exams present far distinct challenges against common image generation tasks:
@@
As shown in Fig. 1, even state-of-the-art text-to-image models, such as GPT-Image-1 (GPT-4o) OpenAI (2025b) and Gemini-2.5-Flash-Image (Nano Banana) Google (2025), excel at producing images with superficially plausible overall structure but often fall short in portraying correct exam details.
Based on this motivation, we introduce GenExam, the first benchmark dedicated to multidisciplinary text-to-image exams. As shown in Fig. 2, GenExam asks the model to accomplish the subject-specific drawing exam with a series of detailed, rigorous and knowledgeable requirements. It includes 1,000 samples of 10 subjects with detailed four-level taxonomy, collected from college-level exams and open-source benchmarks and carefully reviewed by PhD annotators. In addition, GenExam adopts a strict prompt generation process through cooperation of GPT-5 OpenAI (2025a) and humans, with each prompt designed in the style of real exams CollegeBoard (1952); IBO (1968). Given these designs, models are required to seamlessly integrate their understanding, reasoning and generation capabilities to solve the problem of GenExam.
The evaluation of multidisciplinary images also poses challenges, as they prioritize semantic correctness over photorealism or aesthetic quality, different from natural images. Relying solely on a single instruction for MLLM-as-a-judge evaluation Zhang et al. (2025b) makes it difficult to cover all requirements in the prompt. Therefore, we design customized evaluation criteria for each individual sample, similar to the exam marking process. As shown in in Fig. 1, scoring points of each sample are produced through a rigorous review process of GPT-5 and humans, with the ground truth image as the evaluation reference. Leveraging the most advanced commercial MLLM, i.e. GPT-5, we can automatically diagnose each scoring point through visual question answering. In addition to the correctness evaluation, GenExam also includes additional metrics to assess the visual plausibility, namely encompassing spelling, readability, and logical consistency. Based on the above metrics, we calculate two final scores to satisfy different evaluation requirements, with strict and relaxed standards respectively.
Our experimental results show that even state-of-the-art text-to-image models such as GPT-Image-1 OpenAI (2025b) and Gemini-2.5-Flash-Image Google (2025) can only achieve strict scores of less than 15%. Many of the latest T2I and unified multimodal large language models (MLLMs) like FLUX.1 Kontext max Batifol et al. (2025), Qwen-Image Wu et al. (2025a) and BAGEL Deng et al. (2025) even yield almost 0% strict scores, highlighting the great challenge of GenExam. Our experiments also indicate a significant gap between open-source and closed-source models. All these empirical studies suggest a long path ahead for existing models towards general AGI. Overall, our contributions can be summarized in three folds:
@@ 2 Related Work
2.1 Text-to-Image Generation Models
@@
Text-to-image generation (T2I) has been a primary research focus in generative AI in recent years. The dominant approaches are based on diffusion Ho et al. (2020); Dhariwal & Nichol (2021); Rombach et al. (2022); Labs (2024) or autoregressive objects Ramesh et al. (2021); Tian et al. (2024); Sun et al. (2024), which excel at generating images with high aesthetic quality and strong semantic consistency with prompts. Recently, with the development of understanding-only MLLMs Bai et al. (2025b); Luo et al. (2024; 2025a); Wang et al. (2025a; b); Hong et al. (2025), unified MLLMs have emerged as a popular research topic, supporting multimodal understanding and image generation with a single model Xie et al. (2025); Deng et al. (2025); Li et al. (2025); Chen et al. (2025c). The core advantage of unified MLLMs lies in providing powerful multimodal understanding and reasoning capabilities for image generation. The latest state-of-the-art models like GPT-Image-1 OpenAI (2025b) and Gemini-2.5-Flash-Image Google (2025) show strong capabilities in generating plausible natural and aesthetic images. However, as shown in our experiments, their ability on multidisciplinary text-to-image exams is still limited.
2.2 Text-to-Image Generation Benchmarks
Early evaluation of text-to-image generation is mainly based on image similarity metrics, e.g. FID Heusel et al. (2017). These methods often fail to capture the higher-level semantics of the generated images. Current evaluation for T2I focuses primarily on literal prompt-image alignment, with benchmarks like GenEval Ghosh et al. (2023) and metrics like CLIP score Hessel et al. (2021) and VQA score Lin et al. (2024). Inspired by reasoning abilities in pure-text and multimodal tasks Guo et al. (2025a); Yue et al. (2025); Hong et al. (2025); Tang et al. (2024); Team et al. (2025), recent works move towards reasoning-informed generation Niu et al. (2025); Zhao et al. (2025); Meng et al. (2024); Sun et al. (2025); Gong et al. (2025); Zhang et al. (2025a); Wu et al. (2025b), and typically rely on MLLM-as-a-judge Zhang et al. (2025b) for evaluation. However, their focus is mainly on world knowledge or commonsense reasoning, with domains restricted to natural or synthetic images.
@@ 2.2 Text-to-Image Generation Benchmarks
Meanwhile, benchmarks such as MMMG Luo et al. (2025b), OneIG-Bench Chang et al. (2025a), and SridBench Chang et al. (2025b) explore multidisciplinary image generation, but primarily in the form of knowledge-concept illustration, a simple task with loose diagnostic criteria compared to exam-style questions. Existing benchmarks for multidisciplinary images are mainly for understanding tasks Wang et al. (2024b); Yue et al. (2024a); Xia et al. (2024); Zhou et al. (2025); Lin et al. (2025); Phan et al. (2025), but rarely focus on image generation. In contrast, our GenExam targets multidisciplinary text-to-image exams, whose prompts are complex, precise, and diverse and pose strict and explicit constraints on generation. By designing fine-grained scoring points, GenExam provides a more rigorous and accurate evaluation of reasoning-informed generation in multidisciplinary images.
3 GenExam
3.1 Overview
Fig. 2 provides an overview of our GenExam. It contains 1,000 samples that span over 10 subjects: mathematics, physics, chemistry, biology, computer science, geography, economics, music, history, and engineering. Each sample contains a ground truth image, a four-level taxonomy, and several scoring points to judge the semantic correctness of the image.
Scoring Points. One critical challenge in the evaluation of exam-style image generation is that it is difficult to judge the semantic correctness of the generated image. This becomes evident when using MLLM-as-a-judge Zhang et al. (2025b) with a single instruction template, where MLLMs often fail to cover all the requirements specified in the prompt. Examples include the specific structure of organic molecules (e.g., the number of each type of atom, functional groups, chemical bonds), positional relationships in geometric figures, all notes in a musical score, etc. Therefore, inspired by scoring criteria for graph-drawing questions in human exams, we design several scoring points for each prompt, in the form of questions. By answering the questions, we can decide whether the generated image is correct, e.g. “Does the molecule contain exactly 8 carbon atoms?”. This ensures that each sample has a unique evaluation criterion, which improves the accuracy and stability of the evaluation. Each scoring point has a predefined score that sums up to 1, similar to the grading system in exams. The detailed evaluation protocol is given in Sec. 3.3.
Taxonomy. In addition to the top-level subject, we also provide a detailed four-level taxonomy, such as “Mathematics/Plane Geometry/Circles/Tangent” in Fig. 2. The taxonomy annotations are built based on ISCED-F UNESCO (2013) fields, with the aim of improving the academic strictness of GenExam for seamless integration with multidisciplinary research. Detailed four-level taxonomy is provided in Appendix D.
Statistics. In Tab. 1, the prompts in GenExam exhibit an average length of 74.8 words, providing dense and detailed requirements for the generated images, resembling difficult graph-drawing questions in real human exams. Each prompt has an average of 6.9 scoring points to thoroughly assess the generated image from all aspects based on the input prompt. In Fig. 3, we visualize the distributions of the image types, which demonstrate the diversity of images in our benchmark. We also provide a subset of 250 images to facilitate efficient and more affordable evaluation. More statistics are provided in Appendix A.
3.2 Data Curation Pipeline
Data Collection. The data curation pipeline is shown in Fig. 4. To construct a general benchmark for multidisciplinary text-to-image exam, we first consider possible subjects and domains of multidisciplinary images and curate a four-level taxonomy assisted by GPT-5 OpenAI (2025a) and through manual check. We then use these taxonomy to generate keywords related to specific subject knowledge, and collect web images from exams and textbooks based on Google search. Images from existing subject-knowledge-related datasets are also collected: MMMU Yue et al. (2024a), ScienceQA Lu et al. (2022), TextbookQA Kembhavi et al. (2017), MAVIS Zhang et al. (2024), ChEBI Hastings et al. (2016), HLE Phan et al. (2025) and R-Bench Guo et al. (2025b). We obtain approximately 40K images from the data sources with their original text descriptions (e.g. webpage content for web images, lecture notes in TextbookQA, or question and answers in MMMU). We then conduct automatic filtering by utilizing GPT-5 to rate images in terms of text richness, image domain, complexity and subject knowledge density, and set a threshold to remove undesirable images. 6.5K images are left after automatic filtering. Details of automatic filtering and proportion of image sources are in Appendix B.
@@ 3.2 Data Curation Pipeline
Annotating. We employ GPT-5 to generate the initial version of prompts and scoring points. Prompts that require subject knowledge and reasoning are generated, similar to graph-drawing questions in exams. The original text description from the data source is used for reference to improve the precision, bu should be treated carefully since they may contain irrelevant information. Subsequently, scoring points (questions and scores) are generated based on the prompt. We enforce through prompt engineering that the questions cannot be too general (e.g. “Does the image show the correct structure of benzene?”) or too complex (can be split into multiple sub-questions). Scores are based on the difficulty of the questions, where basic questions should have low scores and harder ones have high scores. We provide some few-shot examples to enhance the labeling process. The complete prompts are provided in Appendix F.2.
@@
3.3 Evaluation Protocol
Unlike the evaluation of natural images, multidisciplinary image generation focuses more on correctness than photorealism or aesthetic quality, and a simple flaw can cause significant errors. For instance, if a single atom is drawn mistakenly in a chemical structure or some arrows are depicted in the opposite direction in a diagram, the entire image is incorrect. Similar examples include arrows pointing at the opposite direction, wrongly labeled points, etc. In practice, common failure cases of current T2I models (Fig. 7) are either from semantic consistency with the prompt or plausibility of the image itself (e.g. spelling errors). Therefore, to thoroughly judge the correctness and overall quality of the generated image, we consider two dimensions: semantic correctness and visual plausibility. The evaluation pipeline is shown in Fig. 5.
  1. 1.
    Spelling: Whether the spelling of the text in the image is correct, including notation and equations. This assesses the model’s text rendering capability, which has become the focus in recent T2I models OpenAI (2025b); Wu et al. (2025a).
  2. 2.
  3. 3.
    Readability: Whether all components are clearly readable and identifiable, without incorrectly placed labels and marks, overlap, occlusion, and unlabeled key components. This requires the model to design an appropriate layout and place the components, potentially through reasoning.
Strict and Relaxed Scores. Based on the two dimensions, we first calculate a strict score as the percentage of correct images among all generated images. An image is considered correct if and only if semantic correctness is 1 (i.e. all scoring points are correct) and the scores of spelling, logical consistency and readability are all 2, indicating full satisfaction of all evaluation dimensions. It is only more aligned with our intention for multidisciplinary image generation where a tiny flaw could lead to a technically wrong image, posing a great challenge for T2I models.
However, as shown in Tab. 2, existing methods struggle to generate strictly correct images on GenExam. To highlight the difference between these models, we additionally calculate a relaxed score as the weighted average of the semantic correctness, spelling, logical consistency and readability. The weights are selected based on alignment with human preferences, similar to the process in WiScore Niu et al. (2025), and the values are 0.7, 0.1, 0.1 and 0.1 respectively. We encourage future research to report both strict and relaxed scores for comparison.
Evaluator Model. We utilize MLLM-as-a-judge Zhang et al. (2025b) for the above evaluation with carefully crafted prompts, provided in Appendix F.3. Specifically, we adopt GPT-5 OpenAI (2025a), which shows strong multimodal understanding capabilities in multidisciplinary images. In Appendix Tab. 4, we also compare the robustness of other MLLMs with their API costs and time.
@@ 3.3 Evaluation Protocol
Gemini-2.5-Flash-Image 0.7 43.1 7.1 60.9 4.2 45.3 5.1 72.6 4.5 70.2 4.9 47.4 10.0 65.8 1.3 59.8 1.5 37.0 0.0 57.1 3.9 55.9
Seedream 3.0 0.7 18.6 0.0 21.5 0.8 18.3 0.0 32.2 0.0 38.2 0.0 15.3 0.0 26.5 0.0 12.5 0.0 21.6 0.0 29.2 0.2 23.4
FLUX.1 Kontext max 0.0 23.5 0.0 25.6 0.0 19.2 0.0 38.3 0.0 47.5 0.0 20.9 0.0 28.9 0.0 22.3 0.0 25.4 0.0 33.5 0.0 28.5
@@
Table 2: Strict scores (Str) and relaxed scores (Rel) on GenExam. GPT-Image-1 achieves the highest strict score of 12.1%, followed by Seedream 4.0, Imagen-4-Ultra and Gemini-2.5-Flash-Image. The rest of the models yield nearly 0%. These results highlight the great challenge of multidisciplinary text-to-image exam.
4 Experiments
In this section, we evaluate the performance of representative methods on our GenExam. We also validate the alignment between MLLM-as-a-judge and human evaluation. We use gpt-5-2025-08-07 OpenAI (2025a) as the evaluator model and set the reasoning effort as “low”. More experiments are provided in Appendix C, including results categorized by subject knowledge difficulty and level-2 taxonomy and detailed scores of each dimension.
4.1 Baselines
We select 18 representative models for comparison, including closed-source proprietary models, open-source T2I models and unified MLLMs. We use the default T2I configuration for each of the models for inference.
Closed-source Models: GPT-Image-1 (GPT-4o) OpenAI (2025b), Gemini-2.5-Flash-Image (Nano Banana) Google (2025), Imagen-4-Ultra Deepmind (2025), Seedream 4.0 Seed (2025), Seedream 3.0 Gao et al. (2025), and FLUX.1 Kontext max Batifol et al. (2025). These proprietary models represent the state-of-the-art methods in text-to-image generation.
Open-source T2I Models: Qwen-Image Wu et al. (2025a), FLUX.1 dev Labs (2024), FLUX.1 Krea Lee et al. (2025), HiDream-I1-Full Cai et al. (2025), and Stable Diffusion 3.5 Large Rombach et al. (2022). These models are dedicated to the T2I task.
Open-source Unified MLLMs: Show-o2-7B, Show-o2-1.5B-HQ Xie et al. (2025), BAGEL Deng et al. (2025) (thinking & non-thinking), Janus-Pro Chen et al. (2025c), Emu3 Wang et al. (2024a), BLIP3o-8B Chen et al. (2025a), and BLIP3o-NEXT-GRPO-Text-3B Chen et al. (2025b). These models employ a unified architecture for T2I and multimodal understanding.
4.2 Main Results
The strict and relaxed scores on GenExam are provided in Tab. 2. For strict scores, we observe that all models struggle in multidisciplinary text-to-image exams, with all strict scores lower than 15% and most models nearly 0%, despite their superior capability in general T2I tasks. This demonstrates the great challenge of our benchmark. Specifically, GPT-Image-1 achieves the highest strict score of 12.1%, highlighting its relatively strong ability to integrate multidisciplinary knowledge understanding and reasoning into image generation. Seedream 4.0, Gemini-2.5-Flash-Image and Imagen-4-Ultra are also capable of generating correct images in some cases, with overall strict scores of 7.2%, 6.9% and 3.9%, respectively. Seedream 3.0 and FLUX.1 Kontext max fail in almost all the cases. Open-source models all achieve near-zero strict scores, leaving a huge gap with leading closed-source proprietary models.
The differences between the models are highlighted with relaxed scores. The performance gap between open-source and closed-source models still exists, with GPT-Image-1 being the top-performing method (62.6 relaxed score). Assisted with a multimodal understanding MLLM (Qwen2.5-VL Bai et al. (2025b)) for feature extraction, Qwen-Image achieves the highest relaxed score of 27.5 among open-source methods, showing the necessity of integrating understanding ability into T2I. However, current unified MLLMs like BAGEL and Show-o2 still lag behind dedicated T2I models, suggesting a significant challenge for unified MLLMs to utilize its multidisciplinary knowledge and reasoning ability into image generation.
To gain deeper insights into the strengths and limitations of the models, we provide a comparison of representative models on each evaluation dimension in Fig. 6. Specifically, GPT-Image-1 surpasses other models across all four dimensions, notably with a significant advantage in spelling and logic consistency. Gemini-2.5-Flash-Image, Imagen-4-Ultra and Seedream 4.0 also exhibit leading performance. For open-source models like Qwen-Image and HiDream-I1-Full, their readability is relatively close to the top models, but the other three dimensions still pose great challenges.

### v3

@@ Abstract
Exams are a fundamental test of expert-level intelligence and require integrated understanding, reasoning, and generation. Existing exam-style benchmarks mainly focus on understanding and reasoning tasks, and current generation benchmarks emphasize the illustration of world knowledge and visual concepts, neglecting the evaluation of rigorous drawing exams. We introduce GenExam, the first benchmark for multidisciplinary text-to-image exams, featuring 1,000 samples across 10 subjects with exam-style prompts organized under a four-level taxonomy. Each problem is equipped with ground-truth images and fine-grained scoring points to enable a precise evaluation of semantic correctness and visual plausibility. Experiments on 17 text-to-image and unified models demonstrate the great challenge of GenExam and the huge gap where open-source models consistently lag behind the leading closed-source ones. By framing image generation as an exam, GenExam offers a rigorous assessment of models’ ability to integrate understanding, reasoning, and generation, providing insights for on the path to intelligent generative models. Our benchmark and evaluation code are released at https://github.com/OpenGVLab/GenExam.
Figure 1: Examples of state-of-the-art generative models on the GenExam benchmark. Orange dashed circles indicate scoring points. GenExam contains complex and diverse prompts resembling human exams, and pose great challenge to existing models.
@@ 1 Introduction
Exams are the ultimate test of expert-level intelligence. They are not merely about recalling knowledge points, but serve as a comprehensive assessment of understanding, reasoning, and generation. The ability to solve multidisciplinary exam problems indicates that a model has expert-level intelligence surpassing that of most adults. Therefore, exam-style benchmarks naturally serve as a crucial yardstick to evaluate the progress of artificial general intelligence (AGI) towards domain expertise (Morris et al., 2023; Yue et al., 2024a; Phan et al., 2025).
However, existing multidisciplinary benchmarks are primarily focused on understanding tasks, including pure text benchmarks (Wang et al., 2024b; Zhong et al., 2023) and multimodal benchmarks (Yue et al., 2024a, b), but rarely focus on image generation tasks. In the field of text-to-image generation (T2I), current benchmarks focus more on general world knowledge reasoning (Niu et al., 2025; Sun et al., 2025). Some benchmarks have explored the domain of subject-specific images but are largely limited to concept illustration (Luo et al., 2025; Chang et al., 2025b, a), a relatively simple task with loose diagnostic criteria, similar to “illustrating concepts through image generation” rather than “solving a drawing exam”.
Upon a closer inspection of graph-drawing questions in common multidisciplinary exams, such as AP (CollegeBoard, 1952), A-level (Cambridge, 1951), and IB (IBO, 1968), we find that text-to-image exams present far distinct challenges against common image generation tasks:
@@
As shown in Fig. 1, even state-of-the-art text-to-image models, such as Nano Banana Pro (Google, 2025b) and GPT-Image-1.5 (OpenAI, 2025c) excel at producing images with superficially plausible overall structure but often fall short in portraying correct exam details.
Based on this motivation, we introduce GenExam, the first benchmark dedicated to multidisciplinary text-to-image exams. As shown in Fig. 2, GenExam asks the model to accomplish the subject-specific drawing exam with a series of detailed, rigorous and knowledgeable requirements. It includes 1,000 samples of 10 subjects with detailed four-level taxonomy, collected from college-level exams and open-source benchmarks and carefully reviewed by PhD annotators. In addition, GenExam adopts a strict prompt generation process through cooperation of GPT-5 (OpenAI, 2025a) and humans, with each prompt designed in the style of real exams (CollegeBoard, 1952; IBO, 1968). Given these designs, models are required to seamlessly integrate their understanding, reasoning and generation capabilities to solve the problem of GenExam.
The evaluation of multidisciplinary images also poses challenges, as they prioritize semantic correctness over photorealism or aesthetic quality, different from natural images. Relying solely on a single instruction for MLLM-as-a-judge evaluation (Zhang et al., 2025b) makes it difficult to cover all requirements in the prompt. Therefore, we design customized evaluation criteria for each individual sample, similar to the exam marking process. As shown in in Fig. 1, scoring points of each sample are produced through a rigorous review process of GPT-5 and humans, with the ground truth image as the evaluation reference. Leveraging the most advanced commercial MLLM, i.e. GPT-5, we can automatically diagnose each scoring point through visual question answering. In addition to the correctness evaluation, GenExam also includes additional metrics to assess the visual plausibility, namely encompassing spelling, readability, and logical consistency. Based on the above metrics, we calculate two final scores to satisfy different evaluation requirements, with strict and relaxed standards respectively.
Our experimental results show that although Nano Banana Pro and GPT-Image-1.5 achieves strict scores of 72.7% and 43.2%, respectively, most models can only obtain strict scores less than 15%, and many latest T2I and unified multimodal large language models (MLLMs) like FLUX.2 dev (Labs, 2025), Qwen-Image-2512 (Wu et al., 2025) and BAGEL (Deng et al., 2025) even yield strict scores less than 3%. These results highlight the great challenge of GenExam and the huge gap between closed-source and open-source models, suggesting potential directions for existing models towards AGI. Overall, our contributions can be summarized in three folds:
@@ 2 Related Work
2.1 Generative Models
@@
Text-to-image generation (T2I) has been a primary research focus in generative AI in recent years. The dominant approaches are based on diffusion (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022; Labs, 2024) or autoregressive objects (Ramesh et al., 2021; Tian et al., 2024; Sun et al., 2024), which excel at generating images with high aesthetic quality and strong semantic consistency with prompts. Recently, unified MLLMs have emerged as a popular research topic, which supports multimodal understanding and image generation with a single model (Wang et al., 2024a; Deng et al., 2025; Xie et al., 2025; Li et al., 2025). The core advantage of unified MLLMs lies in providing powerful multimodal understanding and reasoning capabilities for image generation. The latest state-of-the-art models like GPT-Image-1.5 (OpenAI, 2025c) and Nano Banana Pro (Google, 2025b) show strong capabilities in generating plausible natural and aesthetic images. However, as shown in our experiments, their ability on multidisciplinary text-to-image exams is still limited.
2.2 Image Generation Benchmarks
@@
Early T2I evaluation is mainly based on image similarity metrics (Heusel et al., 2017), failing to capture the higher-level semantics. Current evaluation for T2I focuses primarily on literal prompt-image alignment, with benchmarks like GenEval (Ghosh et al., 2023) and metrics like CLIP score (Hessel et al., 2021) and VQA score (Lin et al., 2024). Recent works move towards reasoning-informed generation (Niu et al., 2025; Zhao et al., 2025; Meng et al., 2024; Sun et al., 2025; Zhang et al., 2025a), and typically rely on MLLM-as-a-judge (Zhang et al., 2025b) for evaluation. However, their focus is mainly on world knowledge or commonsense reasoning, with domains restricted to natural or synthetic images. Meanwhile, benchmarks such as MMMG (Luo et al., 2025), OneIG-Bench (Chang et al., 2025a), and SridBench (Chang et al., 2025b) explore multidisciplinary image generation, but primarily in the form of knowledge-concept illustration, a relatively simple task with loose diagnostic criteria compared to exam-style questions. Existing benchmarks for multidisciplinary exams are mainly for understanding tasks (Yue et al., 2024a; Xia et al., 2024; Phan et al., 2025), but rarely focus on image generation. In contrast, our GenExam targets multidisciplinary text-to-image exams, with complex, precise, and diverse prompts and strict and explicit constraints on generation. By designing fine-grained scoring points, GenExam provides a more rigorous evaluation of reasoning-informed generation in multidisciplinary images.
@@
3 GenExam
3.1 Overview
Fig. 2 provides an overview of our GenExam benchmark. It contains 1,000 samples that span over 10 subjects: mathematics, physics, chemistry, biology, computer science, geography, economics, music, history, and engineering. Each sample contains a ground truth image, a four-level taxonomy, and several scoring points.
Scoring Points. One critical challenge in the evaluation of exam-style image generation is that it is difficult to judge the semantic correctness of the generated image. This becomes evident when using MLLM-as-a-judge (Zhang et al., 2025b) with a single instruction template, where MLLMs often fail to cover all the requirements specified in the prompt. Examples include the specific structure of organic molecules (e.g., the number of each type of atom, functional groups, chemical bonds), positional relationships in geometric figures, all notes in a musical score, etc. Therefore, inspired by scoring criteria for graph-drawing questions in human exams, we design several scoring points for each prompt, in the form of questions. By answering the questions, we can decide whether the generated image is correct, e.g. “Does the molecule contain exactly 8 carbon atoms?”. This ensures that each sample has a unique evaluation criterion, which improves the accuracy and stability of the evaluation. Each scoring point has a predefined score that sums up to 1, similar to the grading system in exams. The detailed evaluation protocol is given in Sec. 3.2.
Taxonomy. In addition to the top-level subject, we also provide a four-level taxonomy, such as “Mathematics/Plane Geometry/Circles/Tangent” in Fig. 2. The taxonomy annotations are built based on ISCED-F (UNESCO, 2013) fields, with the aim of improving the academic strictness of GenExam for seamless integration with multidisciplinary research. Detailed taxonomy is provided in Appendix D.
@@ 3.1 Overview
Statistics. In Tab. 1, the prompts in GenExam exhibit an average length of 74.8 words, providing dense and detailed requirements for the generated images, resembling difficult graph-drawing questions in real human exams. Each prompt has an average of 6.9 scoring points to thoroughly assess the generated image from all aspects based on the input prompt. In Fig. 3, we visualize the distributions of the image types, which demonstrate the diversity of images in our benchmark. We also provide a subset to facilitate an efficient evaluation. More statistics are provided in Appendix A.
Data Curation. We first consider possible subjects and domains of multidisciplinary images and curate a four-level taxonomy, and use it to generated keywords for web image collection. Images from existing subject-knowledge-related multimodal understanding datasets are also collected. We then conduct automatic filtering by utilizing GPT-5 to rate images in terms of text richness, image domain, complexity and subject knowledge density, and set a threshold to remove undesirable images. Subsequently, prompts and scoring points are curated through GPT-5 drafting and rigorous human refinement. Detailed data curation pipeline is provided in Sec. B.
@@
Table 2: Strict scores (Str) and relaxed scores (Rel) on GenExam.
3.2 Evaluation Protocol
Unlike the evaluation of natural images, multidisciplinary images focus more on correctness than photorealism or aesthetic quality, and a simple flaw can cause significant errors. For instance, if a single atom is drawn mistakenly in a chemical structure or some arrows are depicted in the opposite direction in a diagram, the entire image is incorrect. In practice, common failure cases of current T2I models (Fig. 6) are either from semantic consistency with the prompt or plausibility of the image itself (e.g. spelling errors). Therefore, to thoroughly judge the correctness and overall quality of the generated image, we consider two dimensions: semantic correctness and visual plausibility. The evaluation pipeline is shown in Fig. 4.
  1. 1.
    Spelling: Whether the spelling of the text in the image is correct, including notation and equations. This assesses the model’s text rendering capability, which has become the focus in recent models (Google, 2025b; Wu et al., 2025).
  2. 2.
  3. 3.
    Readability: Whether all components are clearly readable and identifiable, without incorrectly placed labels and marks, overlap, occlusion, and unlabeled key components. This requires the model to design an appropriate layout and place the components, e.g. through reasoning.
Strict and Relaxed Scores. Based on the two dimensions, we first calculate a strict score as the percentage of correct images among all generated images. An image is considered correct if and only if semantic correctness is 1 (i.e. all scoring points are correct) and the scores of spelling, logical consistency and readability are all 2, indicating full satisfaction of all evaluation dimensions, similar to RISEBench (Zhao et al., 2025). It is only more aligned with multidisciplinary image generation where a tiny flaw could lead to a technically wrong image, posing great challenges for T2I models.
However, as shown in Tab. 2, existing methods struggle to generate strictly correct images on GenExam. To highlight the difference between these models, we additionally calculate a relaxed score as the weighted average of the semantic correctness, spelling, logical consistency and readability. The weights are selected based on alignment with human preferences, similar to the process in WiScore (Niu et al., 2025), and the values are 0.7, 0.1, 0.1 and 0.1 respectively. We encourage future research to report both strict and relaxed scores for comparison.
Evaluator Model. We utilize MLLM-as-a-judge (Zhang et al., 2025b) for evaluation with carefully crafted prompts, provided in Appendix E.3. Specifically, we adopt GPT-5 (OpenAI, 2025a), which shows strong multimodal understanding capabilities in multidisciplinary images. In Tab. 3 and Appendix Tab. 4, we also validate the robustness of Gemini-3-Flash and other MLLMs.